# Data Imputation by Pursuing Better Classification: A Supervised Kernel-Based Method

### v2

@@ 3.2 Stage I: Kernel Matrix Imputation with SVM
$=\boldsymbol{1}\boldsymbol{1}^{\top}+\frac{1}{4\eta}\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{y})\left(\boldsymbol{K}_{k}\odot\mathcal{E}\right)\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{y}).$
@@
Therefore, we need to solve a convex linear constrained quadratic programming problem, which can be efficiently accomplished using existing convex optimization toolboxes, such as the $\mathsf{quadprog}$ function in MATLAB. After obtaining the optimal solution $\boldsymbol{d}^{\ast}$, we can then derive the final approximate solution for the problem (6):
@@
$\min_{\mathcal{E}}$ $\frac{1}{2}\boldsymbol{\alpha}^{\top}\mathrm{diag}(\boldsymbol{y})\left(\boldsymbol{K}_{k}\odot\boldsymbol{K}_{\Delta}\odot\mathcal{E}\right)\mathrm{diag}(\boldsymbol{y})\boldsymbol{\alpha}$ (12)
@@
$\min_{\mathcal{E}}\quad\frac{1}{2}\boldsymbol{\alpha}^{\top}\mathrm{diag}(\boldsymbol{y})\left(\boldsymbol{K}_{k}\odot\boldsymbol{K}_{\Delta}\odot\mathcal{E}\right)\mathrm{diag}(\boldsymbol{y})\boldsymbol{\alpha}+\rho\|\mathcal{E}-\boldsymbol{1}\boldsymbol{1}^{\top}\|_{\mathrm{F}}^{2}.$ (13)
After deriving the closed-form solution for this problem, we utilize a non-expansive operator $\mathcal{P}_{+}(\cdot):\mathcal{S}^{N}\rightarrow\mathcal{S}_{+}^{N}$ to project the solution onto the space of PSD matrices [41, 42, 43]. The operator $\mathcal{P}_{+}(\boldsymbol{X})$ performs eigenvalue decomposition on $\boldsymbol{X}=\boldsymbol{V}\boldsymbol{\Sigma}\boldsymbol{V}^{\top}$ and sets any negative eigenvalues to 0, resulting in the new matrix $\hat{\boldsymbol{X}}=\boldsymbol{V}\boldsymbol{\Sigma}_{+}\boldsymbol{V}^{\top}$. By defining $\Gamma\coloneqq\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{y})\left(\boldsymbol{K}_{k}\odot\boldsymbol{K}_{\Delta}\right)\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{y})$, we obtain the optimal PSD approximation solution for the problem (12):
@@ 3.3 Stage II: Data Imputation with The Given Matrix
By equivalently replacing the objective function and imposing location and range constraints on the imputation results, we obtain the following optimization problem for Stage II:
@@
The essence of this task involves solving a nonlinear system of equations. Additionally, by defining the missing ratio of data features as $m$, the aforementioned system consists of $N(N-1)/2$ equations and $Ndm$ variables. Although the objective function in (16) is non-convex with respect to $\Delta\boldsymbol{X}$, the entries $\{K_{i,j}\}_{i,j=1}^{N}$ still provide abundant supervision information for the data imputation process in scenarios where the number of observed features $(d)$ is significantly smaller than the total number of data $(N)$ and the missing ratio $(m)$ ranges between 0 and 1. To address this, we employ the BCD method to solve the optimization problem mentioned above. In each iteration, we select a specific column from $\Delta\boldsymbol{X}$ as a variable while keeping the remaining columns constant. We update the variable iteratively until the maximum number of iterations is reached or the change in the objective function is smaller than the predefined threshold. The accuracy of this solution is further validated by the experimental results presented later in this paper. The complete two-stage data imputation framework is summarized in Algorithm 1.
@@
6:     Solve problem (10) to obtain $d^{\ast}$ and update $\boldsymbol{K}_{\Delta}^{(t_{1})}$ by (11).
7:     Update $\mathcal{E}^{(t_{1})}$ by (14).
@@
15:     for $i=1$ to $N$ do
16:         Fix $\{\Delta\boldsymbol{x}_{j}^{(t_{2}-1)}:j\neq i\}$ and update $\Delta\boldsymbol{x}_{i}^{(t_{2})}$ in (16).
@@ 4.1 Experimental Settings
Data sets and preprocessing. We chose six real-world data sets from libsvm [45] and UCI [46], namely australian, german, heart, pima (a.k.a. diabetes), cylinder, and horse. The details of these data sets are shown in Table 1. The data sets were then divided into three subsets: a training set, a complete holdout set for parameter selection, and a complete test set for evaluating and comparing the final results of the algorithms. The split was performed in a 4:3:3 ratio. Additionally, for the training set of the first four complete data sets, we constructed the missing data by randomly removing a given percentage of the features to evaluate the performance of various algorithms under different levels of missingness. The missing ratio of a data set was defined as
@@
Table 1: Details of the six libsvm and UCI data sets.
@@
[…]
@@
[…]
@@
  • Ours: To avoid excessive parameter tuning, we utilized the parameters $C_{\mathrm{MI}}$ and $\gamma_{\mathrm{MI}}$ from MI and further selected $C\in\{C_{\mathrm{MI}}\cdot 2^{i},\ i=-1,0,1\}$ and $\gamma\in\{\gamma_{\mathrm{MI}}\cdot 2^{i},\ i=-1,0,1\}$. Additionally, we empirically set $\eta=\|\boldsymbol{\alpha}_{\mathrm{MI}}\|_{2}$ and $\rho=5C/m$ during each iteration.
@@ 4.2.1 Results of Data Imputation from Given Kernel Matrix
Ours 0.756 $\pm$ 0.026 0.759 $\pm$ 0.022 0.744 $\pm$ 0.027 $\bullet$ 0.709 $\pm$ 0.037 $\bullet$
@@ 4.2.2 Classification Results on Real-World Data Sets
In this section, we compared different data imputation approaches using six real-world data sets. Since all imputation algorithms ultimately serve subsequent tasks, we evaluated their performance by calculating the mean and variance of their classification accuracy on the test data.
We first mimicked the missing completely at random (MCAR) mechanism by randomly removing some values on the four complete data sets, and the results were presented in Table 3. For situations with a relatively low missing data ratio, such as $m=20\%$, the differences between the methods are not significant. Even using a simple method like MI can achieve decent predictive performance. When $m=40\%$, our proposed method achieves the highest average accuracy and the lowest standard deviation on the australian, heart, and pima data sets, demonstrating the superior stability of our method. On the german data set, our algorithm’s performance is second only to genRBF. When dealing with high missing data ratios, the prediction accuracy of the GEOM, KARMA, and genRBF methods fluctuates across different data sets. However, our method demonstrates even more significant advancements in such scenarios. For instance, at $m=60\%$, our approach outperforms the second-ranking method by an additional accuracy improvement of approximately $2.6\%$. This significant improvement has also been confirmed through a paired t-test conducted at the 5% significance level. This performance is further reflected in the case of $m=80\%$, where our algorithm achieves precise and stable classification tasks.
@@
Data Sets $m$ Methods Accuracy
Cylinder $\approx 10\%$ MI 0.648 $\pm$ 0.014
GEOM 0.620 $\pm$ 0.021
KARMA 0.678 $\pm$ 0.018
genRBF 0.620 $\pm$ 0.021
Ours 0.672 $\pm$ 0.020
Horse $\approx 30\%$ MI 0.844 $\pm$ 0.027
GEOM 0.848 $\pm$ 0.021
KARMA 0.842 $\pm$ 0.020
genRBF 0.836 $\pm$ 0.028
Ours 0.862 $\pm$ 0.029

### v3

@@ 3.2 Stage I: Kernel Matrix Imputation with SVM
$=\boldsymbol{1}\boldsymbol{1}^{\top}+\frac{1}{4\eta}\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{y})\left(\boldsymbol{K}_{o}\odot\mathcal{E}\right)\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{y}).$
@@
Therefore, we need to solve a convex linear constrained quadratic programming problem, which can be efficiently accomplished using existing convex optimization toolboxes, such as the quadprog function in MATLAB. After obtaining the optimal solution $\boldsymbol{d}^{\ast}$, we can then derive the final approximate solution for the problem (6):
@@
$\min_{\mathcal{E}}$ $\frac{1}{2}\boldsymbol{\alpha}^{\top}\mathrm{diag}(\boldsymbol{y})\left(\boldsymbol{K}_{o}\odot\boldsymbol{K}_{\Delta}\odot\mathcal{E}\right)\mathrm{diag}(\boldsymbol{y})\boldsymbol{\alpha}$ (12)
@@
$\min_{\mathcal{E}}\quad\frac{1}{2}\boldsymbol{\alpha}^{\top}\mathrm{diag}(\boldsymbol{y})\left(\boldsymbol{K}_{o}\odot\boldsymbol{K}_{\Delta}\odot\mathcal{E}\right)\mathrm{diag}(\boldsymbol{y})\boldsymbol{\alpha}+\rho\|\mathcal{E}-\boldsymbol{1}\boldsymbol{1}^{\top}\|_{\mathrm{F}}^{2}.$ (13)
After deriving the closed-form solution for this problem, we utilize a non-expansive operator $\mathcal{P}_{+}(\cdot):\mathcal{S}^{N}\rightarrow\mathcal{S}_{+}^{N}$ to project the solution onto the space of PSD matrices [43, 44]. The operator $\mathcal{P}_{+}(\boldsymbol{X})$ performs eigenvalue decomposition on $\boldsymbol{X}=\boldsymbol{V}\boldsymbol{\Sigma}\boldsymbol{V}^{\top}$ and sets any negative eigenvalues to 0, resulting in the new matrix $\hat{\boldsymbol{X}}=\boldsymbol{V}\boldsymbol{\Sigma}_{+}\boldsymbol{V}^{\top}$. By defining $\Gamma\coloneqq\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{y})\left(\boldsymbol{K}_{o}\odot\boldsymbol{K}_{\Delta}\right)\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{y})$, we obtain the optimal PSD approximation solution for the problem (12):
@@
Additionally, for the non-trivial min-max problem with box and PSD constraints presented in problem (5), although it can be solved through the above three alternating optimization steps, its convergence remains unclear. In B, we provide a comprehensive review of the entire procedure, including detailed definitions of all notations, and then, based on several lemmas, present the following theorem concerning the norm difference of the $\boldsymbol{K}_{\Delta}$ parameters during iterations.
Theorem 1.
When $t$ denotes the iteration index, the difference in $\boldsymbol{K}_{\Delta}$ between two consecutive updates can be bounded by
$\left\|\boldsymbol{K}_{\Delta}^{(t+1)}-\boldsymbol{K}_{\Delta}^{(t)}\right\|_{\mathrm{F}}\leq L\left\|\boldsymbol{K}_{\Delta}^{(t)}-\boldsymbol{K}_{\Delta}^{(t-1)}\right\|_{\mathrm{F}}+R^{(t)},$ (15)
where $L=\frac{1}{4\eta}N^{2}C^{4}\kappa^{2}$ and $R^{(t)}=\frac{1}{2\eta}\nu^{(t)}N^{2}C^{3}\kappa^{2}b_{\mathrm{max}}\left(1+C\kappa b_{\mathrm{max}}+\frac{NC^{3}\kappa\lambda b_{\mathrm{max}}^{2}}{4\rho}\right)+\frac{1}{2\eta}\nu^{(t)}NC\kappa$
$\left(1+\frac{NC^{2}\kappa b_{\mathrm{max}}}{4\rho}\right)\left(1+C\kappa b_{\mathrm{max}}+\frac{NC^{3}\kappa\lambda b_{\mathrm{max}}^{2}}{4\rho}\right)$.
Remark 1.
In our algorithm, $\eta$, $\rho$, and $\nu^{(t)}$ control the optimization step sizes for variables $\boldsymbol{K}_{\Delta}$, $\mathcal{E}$, and $\boldsymbol{\alpha}$ in each iteration, respectively. When we set $\eta>\frac{N^{2}C^{4}\kappa^{2}}{4}$, $\rho=\mathcal{O}\left(NC^{3}\kappa\lambda b_{\mathrm{max}}^{2}\right)$, and ensure that $\nu^{(t)}=\mathcal{O}(1/t)$ holds, it will follows that $L<1$ and $R^{(t)}\rightarrow 0$ as the iteration index $t$ increases. Since the actual upper bound may be smaller, we thus establish a sufficient condition that guarantees the convergence of the aforementioned alternating optimization.
@@ 3.3 Stage II: Data Imputation with the Given Matrix
By replacing the objective function and imposing location and range constraints on the imputation results, we obtain the following optimization problem for Stage II:
@@
The essence of this task involves solving a nonlinear system of equations. By defining the missing ratio of data features as $m$, the aforementioned system consists of N(N1)/2N(N-1)/2 equations and $Ndm$ variables. Although the objective function in problem (17) is non-convex to $\Delta\boldsymbol{X}$, the entries $\{K_{i,j}\}_{i,j=1}^{N}$ still provide abundant supervision information for the data imputation process in scenarios where the number of observed features $(d)$ is significantly smaller than the total number of data $(N)$ and the missing ratio $(m)$ ranges between 0 and 1. To address this, we employ the BCD method to solve the optimization problem mentioned above. In each iteration, we select a specific column from $\Delta\boldsymbol{X}$ as a variable while keeping the remaining columns constant. We update the variable iteratively until the maximum number of iterations is reached or the change in the objective function is smaller than the predefined threshold. The accuracy of this solution is further validated by the experimental results presented later in this paper. The complete two-stage data imputation framework is summarized in Algorithm 1.
In addition, we are equally concerned about the time complexity of the algorithm. First, we analyze Stage I. When computing the preliminary results of $\hat{\boldsymbol{K}}_{\Delta}$, $\hat{\mathcal{E}}$, and $\hat{\boldsymbol{\alpha}}$ (excluding the projection operation), the relevant Hadamard products and matrix-vector multiplications require a complexity of $\mathcal{O}(N^{2})$. As for $\boldsymbol{K}_{\Delta}$ and $\mathcal{E}$, their subsequent PSD projection operators and quadratic programming demand a higher complexity of $\mathcal{O}(N^{3})$. Next, for Stage II, the complexity required for single-column updates $\Delta\boldsymbol{x}_{i}$ is $\mathcal{O}(Nd)$, so traversing the complete $\Delta\boldsymbol{X}$ once would require $\mathcal{O}(N^{2}d)$. Therefore, in the scenario we are concerned with (where $N\gg d$), the computational cost of Algorithm 1 in each iteration will be dominated by $\mathcal{O}(N^{3})$, where the empirical number of iterations can be referenced from the results in Figure 6.
@@
6:  Solve problem (10) to obtain $d^{\ast}$ and update $\boldsymbol{K}_{\Delta}^{(t_{1})}$ by equation (11).
7:  Update $\mathcal{E}^{(t_{1})}$ by equation (14).
@@
15:  for $i=1$ to NN do
16:   Fix $\{\Delta\boldsymbol{x}_{j}^{(t_{2}-1)}:j\neq i\}$ and update $\Delta\boldsymbol{x}_{i}^{(t_{2})}$ in problem (17).
@@ 4.1 Experimental Settings
Datasets and preprocessing. We chose seven real-world datasets from libsvm [46] and UCI [47], namely australian, german, heart, pima (a.k.a. diabetes), wine, cylinder, and horse. The details of these datasets are shown in Table 1. The datasets were then divided into three subsets: a training set, a complete holdout set for parameter selection, and a complete test set for evaluating and comparing the final results of the algorithms. The split was performed in a 4:3:3 ratio. Additionally, for the training set of the first five complete datasets, we constructed the missing data by randomly removing a given percentage of the features to evaluate the performance of various algorithms under different levels of missingness. The missing ratio of a dataset was defined as
@@
Table 1: Details of the seven libsvm and UCI datasets.
@@
  • 1.
    MI: The missing values are imputed by replacing them with the mean value of the corresponding features in the training set. We chose the penalty parameter $C\in\{2^{-5},\ 2^{-4},\ \ldots\ ,2^{5}\}$ and the bandwidth in the Gaussian kernel $\gamma\in\{2^{-5},\ 2^{-4},\ \ldots\ ,2^{5}\}$ using the holdout set.
  • 2.
@@
  • 3.
    KARMA [31]: This method proposed a new kernel function for the missing data $k_{\beta}(\cdot,\cdot):(\mathbb{R}\cup\{\ast\})^{d}\times(\mathbb{R}\cup\{\ast\})^{d}\rightarrow\mathbb{R}$. We chose $C\in\{2^{-5},\ 2^{-4},\ \ldots\ ,2^{5}\}$ and $\beta\in\{1,2,3,4\}$ using the holdout set.
  • 4.
@@
  • 5.
    NewImp [29]: This method employs the Wasserstein gradient flow framework and performs data imputation based on a diffusion model. For the hyper-parameters $h$, $\mathrm{HU}_{\mathrm{score}}$, $\lambda$, $\eta$, we follow the grid range described in their paper. Additionally, for SVM, $C$ and $\gamma$ were chosen the same as in MI.
  • 6.
    Ours: To avoid excessive parameter tuning, we utilized $C_{\mathrm{MI}}$ and $\gamma_{\mathrm{MI}}$ from MI and further selected $C\in\{C_{\mathrm{MI}}\cdot 2^{i},\ i=-1,0,1\}$ and $\gamma\in\{\gamma_{\mathrm{MI}}\cdot 2^{i},\ i=-1,0,1\}$. Additionally, we empirically set $\eta=\|\boldsymbol{\alpha}_{\mathrm{MI}}\|_{2}$ and $\rho=5C/m$ during each iteration.
@@ 4.2.1 Results of Data Imputation from Given Kernel Matrix
4.2.2 Classification Results on Artificial datasets
To validate the robustness of our method first, we conducted evaluations using artificially generated datasets with controlled characteristics. These datasets were designed to systematically investigate the influence of four key factors: dimensionality, noise intensity, extreme outliers, and non-uniform distributions. Specifically, we generated two classes of data as follows: $\boldsymbol{X}_{1}\sim\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1})$ and $\boldsymbol{X}_{2}\sim\mathcal{N}(\boldsymbol{\mu}_{2},\boldsymbol{\Sigma}_{2})$, where $\boldsymbol{\mu}_{1}=-\boldsymbol{1}_{d}$, $\boldsymbol{\mu}_{2}=\boldsymbol{1}_{d}$ and $\boldsymbol{\Sigma}_{1}=\boldsymbol{\Sigma}_{2}=\mathrm{diag}(1+\mathrm{non\_uniformity}\cdot\boldsymbol{u}_{d})$, with $\boldsymbol{u}_{d}\sim\mathrm{Uniform}(\boldsymbol{0}_{d},\boldsymbol{1}_{d})$. To simulate noise intensity, Gaussian noise was added to each feature. The noisy dataset is represented as $\boldsymbol{X}_{\mathrm{noisy}}=\boldsymbol{X}+\mathrm{noise\_intensity}\cdot\mathcal{N}(0,1)$, where $\boldsymbol{X}$ represents the combined dataset from both classes. The effect of extreme outliers was simulated by replacing a fraction of the samples with outlier values, generated as $\boldsymbol{X}_{\mathrm{outliers}}=\boldsymbol{X}+10\cdot(\boldsymbol{U}-0.5)$, where $\boldsymbol{U}_{i,j}$ is drawn from a uniform distribution $\mathrm{Uniform}(0,1)$. The effects of these parameters on imputation and subsequent classification tasks are summarized in Figure 3. For all experiments, the default parameter values were set as $N=400$, $d=30$, $\mathrm{noise\_intensity}=0.2$, $\mathrm{outlier\_fraction}=0.1$, $\mathrm{non\_uniformity}=0.5$, and $C=\gamma=1$ for the SVM classifier. Our results show that the proposed method demonstrates strong robustness across different scenarios. As expected, increasing the noise intensity, fraction of outliers, and non-uniformity slightly reduces the accuracy and increases the variance. However, the overall mean accuracy remains relatively stable.
Figure 3: The performance of the proposed method (mean $\pm$ std) on artificial datasets with varying dimensions ($d$), noise intensity, outlier fraction, and non-uniformity.
@@ Methods
NewImp 0.762 $\pm$ 0.012 0.749 $\pm$ 0.023 0.729 $\pm$ 0.022 0.689 $\pm$ 0.030
Ours 0.756 $\pm$ 0.026 0.759 $\pm$ 0.022 0.744 $\pm$ 0.027 $\bullet$ 0.709 $\pm$ 0.037
@@ 4.2.3 Classification Results on Real-World datasets
In this section, we compared different data imputation approaches using seven real-world datasets. Since all imputation algorithms ultimately serve subsequent tasks, we evaluated their performance by calculating the mean and variance of their classification accuracy on the test data.
We first mimicked the missing completely at random (MCAR) mechanism by randomly removing some values on the four complete datasets, and the results were presented in Table 3. For situations with a relatively low missing data ratio, such as $m=20\%$, the differences between the methods are not significant. Even using a simple method like MI can achieve decent predictive performance. When $m=40\%$, our proposed method achieves the highest average accuracy and the lowest standard deviation on the australian, heart, and pima datasets, demonstrating the superior stability of our method. On the german dataset, our algorithm’s performance is only slightly lower than genRBF and NewImp. When dealing with high missing data ratios, the prediction accuracy of the GEOM, KARMA, and genRBF methods fluctuates across different datasets. However, our method demonstrates even more significant advancements in such scenarios. For instance, at $m=60\%$, our approach outperforms the second-ranking method by an additional accuracy improvement of approximately $1.7\%$. This significant improvement has also been confirmed through a paired t-test conducted at the 5% significance level. This performance is further reflected in the case of $m=80\%$, where our algorithm achieves precise and stable classification tasks.
@@
Methods Datasets
Cylinder $(m\approx 10\%)$ Horse $(m\approx 30\%)$
MI 0.648 $\pm$ 0.014 0.844 $\pm$ 0.027
GEOM 0.620 $\pm$ 0.021 0.848 $\pm$ 0.021
KARMA 0.678 $\pm$ 0.018 0.842 $\pm$ 0.020
GenRBF 0.620 $\pm$ 0.021 0.836 $\pm$ 0.028
NewImp 0.657 $\pm$ 0.013 0.854 $\pm$ 0.021
Ours 0.672 $\pm$ 0.020 0.862 $\pm$ 0.029
@@
In addition, some studies have evaluated imputation performance by training multi-layer perceptrons (MLPs) on imputed data [18, 48]. To test beyond kernel-based methods and SVM, we followed the experimental setup in [18] and selected MI, NewImp, and our method for comparison (as GEOM, KARMA, and genRBF are only applicable to kernel machines). For the previously tested datasets australian, german, heart, and pima, we set the missing rate to $m=50\%$ and reported the results of imputation followed by MLP in Table 5. It can be observed that our method still achieves the best classification accuracy in the vast majority of cases and produces more stable results in multiple experiments on the australian and pima datasets. This demonstrates that our label-guided imputation method not only performs well with SVM but also extends its advantages to other classifiers.
Table 5: Comparison of classification accuracy (mean $\pm$ std) of different data imputation methods + MLP. The best performance is highlighted in bold.
Methods /
datasets
$m=50\%$
Australian German Heart Pima
MI 0.854 $\pm$ 0.025 0.710 $\pm$ 0.026 0.792 $\pm$ 0.031 0.721 $\pm$ 0.022
NewImp 0.857 $\pm$ 0.020 0.732 $\pm$ 0.026 0.796 $\pm$ 0.028 0.729 $\pm$ 0.033
Ours 0.866 $\pm$ 0.017 0.723 $\pm$ 0.033 0.806 $\pm$ 0.038 0.740 $\pm$ 0.018
Figure 4: Classification accuracy (mean $\pm$ std) on the wine dataset across different splits (denoted as $s$) for our method with two baselines (MI and NewImp).
Finally, in the previous experiments, the scale of the datasets remained small to moderate. Since our algorithm focuses on leveraging pairwise information between data points to guide the imputation process, handling an $N\times N$ matrix is unavoidable, which makes it challenging to directly scale to large datasets. To address this limitation, we split the dataset into subsets and applied our algorithm to each subset to reduce computational costs. The imputation results from each subset were then combined. Subsequently, we compared the imputation results obtained by MI, NewImp, and our method under different dataset splits and applied them to downstream MLP classification tasks to evaluate performance. The results are presented in Figure 4. As shown in the results, the accuracy of our method for downstream tasks does exhibit a slight decline as the number of subsets increases. However, when the number of splits is small (e.g., 2 or 5), it still performs comparably to NewImp.
Table 6: Comparison of inter-class distance (ICD), Fisher discriminant ratio (FDR), Calinski-Harabasz index (CHI), and Davies-Bouldin index (DBI) values of different imputation methods on different datasets. $(\uparrow)$ means that we expect the value to be larger, $(\downarrow)$ vice versa.
datasets
Methods
Metrics
ICD()($\uparrow$) FDR()($\uparrow$) CHI()($\uparrow$) DBI()($\downarrow$)
Australian
MI 0.4329 0.1608 21.655 3.416
NewImp 0.4363 0.1592 21.411 3.407
Ours 0.4398 0.1666 22.416 3.365
German
MI 0.2603 0.0263 4.480 8.537
NewImp 0.2762 0.0280 4.781 8.243
Ours 0.3873 0.0647 10.989 6.658
Heart
MI 0.4455 0.1481 7.738 3.619
NewImp 0.5196 0.1963 10.280 3.207
Ours 0.5730 0.2745 14.356 3.062
Pima
MI 0.1112 0.0592 8.597 5.249
NewImp 0.1022 0.0510 7.426 6.037
Ours 0.1180 0.0660 9.495 5.042
4.3 Structural Differences of Different Data Imputation Methods
In this section, we explore the structural differences between various imputation methods from both numerical and visual perspectives. Numerically, we employ four clustering metrics: inter-class distance (ICD), Fisher discriminant ratio (FDR), Calinski-Harabasz index (CHI), and Davies-Bouldin index (DBI), to measure whether the imputed data tends to overlap or remains relatively well-separated, which could affect subsequent classification tasks. The definitions of these metrics are provided in C, and the results are summarized in Table 6. From the results, it can be observed that our method achieves larger inter-class distances and better distinguishability in multiple scenarios, with more significant differences from MI and NewImp on the german and heart datasets. These differences can be more intuitively observed in Figure 5, where we present the results on the heart dataset and perform principal component analysis (PCA) dimensionality reduction visualizations for each method. Specifically, to ensure fairness, we used the PCA projection matrix obtained from the complete data to project the imputed data of MI, NewImp, and our method. Clearly, the samples imputed by MI and NewImp remain overlapped, whereas the proposed method demonstrates more distinct separation between different classes in the imputed two-dimensional subspace. This improved separation is expected to be more beneficial for subsequent classification tasks.
(a) MI
(b) NewImp
(c) Ours
Figure 5: PCA visualization of different imputation methods on heart dataset.